# Cross-View UAV Geo-Localization with Precision-Focused Efficient Design: A Hierarchical Distillation Approach with Multi-view Refinement

### v1

@@ IV-B Comparison with State-of-the-Art Methods
We conduct a comprehensive evaluation of our proposed Precision-Focused Efficient Distillation (PFED) framework against a wide array of recent state-of-the-art methods. As presented in Table I, the comparison focuses on both retrieval accuracy (Recall@1, AP) and model efficiency (Parameters, FLOPs, FPS). Our results demonstrate that PFED not only establishes a new state-of-the-art in accuracy but also fundamentally redefines the efficiency frontier for this task.
Superior Efficiency. Our core model, PFED, exhibits high computational efficiency. With only 28.57M parameters and an exceptionally low 4.45G FLOPs, it is by far the most lightweight and computationally inexpensive model among all compared methods. To put this in perspective, compared to the recent efficient model MEAN [4], PFED reduces FLOPs by a remarkable factor of 5.8 and parameters by 21.2%. This optimization directly translates to a massive 3.3 increase in FPS, reaching 1022 images/second, a critical capability for real-time deployment on resource-constrained platforms.
State-of-the-Art Accuracy. Building upon this highly efficient foundation, our full model, PFED†, incorporates the Multi-view Refinement Module (MRM) as a post-processing step achieves an impressive 97.15% Recall@1 and 97.50% AP in the primary DroneSatellite retrieval task. This represents a substantial absolute improvement of 3.30% in R@1 over our efficient baseline and significantly surpasses the previous best-performing method, DAC [27], all while operating with over 20 fewer FLOPs (4.45G vs. 90.24G).
In summary, the results in Table I unequivocally show that PFED successfully achieves a trade-off between accuracy and efficiency in cross-view geo-localization, making it a highly practical and effective solution for real-world applications.
@@ IV-C Unsupervised Domain Adaptation Results
As demonstrated in Table II, our proposed PFED exhibits exceptional generalization in the DroneSatellite task. While MEAN [4] shows strong performance at 150m, our model consistently outperforms all state-of-the-art methods at 200m, 250m, and 300m. Notably, at the 300m range, PFED achieves a top-ranking R@1 score of 97.87, surpassing the next-best competitor by a significant 3.24 percentage points. It is crucial to highlight that PFED utilizes only 4.45G FLOPs, compared to the 26.18G FLOPs required by MEAN. This demonstrates that our model’s feature representation is more robust to severe scale variations while being remarkably lightweight.
The superior generalization of PFED is more pronounced in the SatelliteDrone task, as shown in Table III. In this scenario, our model unequivocally achieves the best performance across all evaluation altitudes. For instance, it surpasses the strong baseline DAC [27] by 3.75 percentage points in R@1 at the 250m altitude. This state-of-the-art accuracy is delivered with a model that is substantially more efficient, requiring only 28.57M parameters compared to DAC’s 96.50M.
IV-D Multi-weather UAV imagery degradation Results
As shown in Figure 5, to evaluate the model’s robustness in practical scenarios, we conducted extensive experiments on the University-1652 dataset under various environmental degradations, including fog, rain, darkness, and wind. As presented in Table IV, these tests assess the model’s ability to maintain high accuracy even when the visual quality of UAV imagery is severely compromised.
In the DroneSatellite retrieval task, our proposed PFED demonstrates remarkable resilience and establishes a new state-of-the-art across all tested conditions. From normal weather to the most severe degradations like darkness and combined fog with rain, PFED consistently achieves the highest Recall@1 and AP scores. For instance, under dark conditions, it outperforms the next-best method, MEAN [4], by a substantial margin of 5.37 percentage points in R@1 (93.27% vs. 87.90%). This superior performance is even more impressive given its efficiency; it delivers these results while requiring 5.8 times fewer FLOPs than MEAN.
In the more challenging SatelliteDrone task, PFED continues to show highly competitive performance. In the most complex scenarios such as ”Fog+Rain” and ”Wind”, PFED takes the lead in both metrics, achieving the best R@1 and AP scores. This proves that even when satellite images must be matched against heavily distorted drone views, our model’s feature representation remains robust and effective.
@@ IV-D Multi-weather UAV imagery degradation Results
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 70.05 | 74.93 | 80.68 | 83.90 | 87.35 | 89.72 | 90.03 | 91.91 |
| DAC[27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 76.65 | 80.56 | 86.45 | 89.00 | 92.95 | 94.18 | 94.53 | 95.45 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 81.73 | 87.72 | 89.00 | 91.00 | 92.13 | 93.60 | 94.63 | 95.76 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 79.85 | 83.32 | 91.10 | 92.79 | 95.57 | 96.45 | 97.87 | 98.32 |
@@
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 83.75 | 73.83 | 91.25 | 83.42 | 93.75 | 89.07 | 93.75 | 90.66 |
| DAC[27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 87.50 | 79.87 | 96.25 | 88.98 | 95.00 | 92.81 | 96.25 | 94.00 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 91.25 | 81.50 | 96.25 | 89.55 | 95.00 | 92.36 | 96.25 | 94.32 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 92.50 | 83.81 | 97.50 | 91.75 | 98.75 | 94.59 | 97.50 | 96.04 |
@@
| PFED (Ours) | 4.45 G | 93.27 | 94.34 | 95.44 | 89.95 |
@@ IV-E Anti-offset Generalization Results
In real-world UAV-view geo-localization, the captured query image is often not perfectly centered or aligned with its corresponding satellite-view gallery image. This spatial misalignment can be caused by variations in camera angle or differences in viewpoint. A robust model must be able to generalize well despite such spatial offsets.
@@
The comprehensive results are presented in Table V. The findings demonstrate the superior robustness of our proposed PFED model. For example, under the most severe shift of pixels, where other models experience a significant performance collapse, PFED maintains an exceptional Rank-1 accuracy of . This represents a massive improvement of +16.92% over the second-best method, SDPL.
It is crucial to note that while the performance of all models degrades as the offset magnitude increases, PFED’s degradation is significantly more graceful. This highlights its ability to learn features that are inherently invariant to spatial translations. This anti-offset generalization capability, coupled with its significantly lower computational cost (4.45 GFLOPs) compared to competitors like FSRA (98.05 GFLOPs) and SDPL (69.71 GFLOPs), underscores the practical value and efficiency of our method for real-world deployment.
| Padding Pixel | FSRA (98.05 G) | LPN (65.39 G) | SDPL (69.71 G) | PFED (4.45 G) | ||||
@@
| HD-CVGL | MRM | DroneSatellite | SatelliteDrone | ||||
@@ IV-F Ablation Studies
Our PFED framework is architected training-time optimization and lightweight inference-time post-process. To validate the specific contributions of each phase, we conducted extensive ablation studies, with the results summarized in Table VI.
@@
Effectiveness of the Training-Phase Framework. The core objective of our training strategy is to distill rich, view-invariant knowledge into a single, efficient network, avoiding the need for multi-branch architectures during inference. Our analysis starts with a baseline model (row 1) achieving 86.44% Recall@1. By introducing Self-Distillation (SD), performance improves significantly to 91.46%, which ensures that the model can fully leverage its own multi-level feature hierarchy for more robust representation learning. Following this, integrating the UAFA module further advances the Recall@1 to 91.93%. Finally, Cross-Distillation (CD) elevates performance to 93.87%, completing the HD-CVGL framework. This step-wise improvement validates our training-phase design, which produces a powerful feature extractor ready for efficient inference. As visualized in Figure 7 and Figure 8, our PFED model learns a more discriminative feature distribution compared to the baseline. The features produced by PFED exhibit smaller intra-class variance and larger inter-class distance, indicating a well-structured feature space.
@@
Effectiveness of the Inference-Phase Module. After establishing a strong student model through our training strategy, we evaluate the contribution of the MRM. The final row of Table VI shows the impact of applying MRM to the descriptors generated by our fully trained model. The result is a remarkable jump in performance to 97.15% Recall@1. This +3.28% gain over the already powerful base model demonstrates that by refining the feature set at inference time, MRM significantly enhances localization precision, fulfilling the ”Precision-Focused” aspect of our PFED paradigm without altering the underlying network architecture. Figure 9 illustrates the performance of different selected views. The model achieves its peak performance with 40 views. Even at its lowest setting (20 views), our approach still significantly outperforms the baseline without the MRM module. As demonstrated by the top-5 retrieval results on the University-1652 and SUES-200 datasets (Figure 10 and Figure 11), PFED exhibits robust cross-view matching capabilities.
@@
Edge Device Deployment. To demonstrate the practical applicability of our approach, as shown in Figure 12, we evaluate PFED on two representative edge platforms: NVIDIA Jetson TX2 and AGX Orin. As shown in Table VII, our method achieves exceptional efficiency with only 4.45 GFLOPs, representing a 95.1% reduction compared to DAC. This dramatic reduction translates directly to superior real-time performance: PFED achieves 27.1 FPS on the resource-constrained TX2 (7.7 faster than DAC’s 3.5 FPS) and an impressive 251.5 FPS on AGX Orin (6.1 faster than DAC). Even compared to the efficient MEAN method, our approach demonstrates 2.9 and 2.8 speedups on TX2 and Orin respectively.
@@ IV-G Limitations and Future Works
Despite its promising performance, PFED has several limitations that guide our future work. The current model, while excelling at DroneSatellite task exhibits an asymmetric performance due to its focus on the multi-view refinement of drone features. Future efforts will explore the intra-relationships within the satellite data to build a more generalized cross-view feature space. Additionally, PFED’s lightweight design faces challenges with extreme low-altitude distortions, which could be mitigated by developing adaptive, scale-aware mechanisms or pre-transformation techniques. Finally, to bridge the gap between efficient image-level retrieval and the need for precise localization, our ultimate goal is to evolve beyond static images. We plan to leverage multiview sequence data and develop an end-to-end framework that integrates coarse scene identification with fine-grained pose estimation, enabling robust and accurate geo-localization within large-scale satellite imagery.
@@ V Conclusion
In this paper, we introduced Precision-Focused Efficient Design (PFED), a framework that resolves the critical accuracy-efficiency trade-off in cross-view geo-localization for UAVs. Our core strategy shifts computational complexity to the training stage, enabling superior performance on resource-constrained devices. We achieve this through two innovations: 1) A comprehensive Hierarchical Distillation (HD-CVGL) strategy during training, which incorporates our Uncertainty-Aware Prediction Alignment (UAPA) to robustly handle data imbalance and domain discrepancies, producing a highly discriminative yet compact student network without any inference overhead. 2) A lightweight Multi-view Refinement Module (MRM) at inference, which uses mutual information to select and fuse the most informative views, boosting accuracy while minimizing feature matching cost. Experiments show PFED sets a new state-of-the-art, achieving 97.15% Recall@1 on University-1652 while being over more FLOPs-efficient and faster than previous methods. Crucially, its 251.5 FPS on an AGX Orin validates its practicality for real-time UAV-based geo-localization in GNSS-denied environments.

### v2

@@ IV-B Comparison with State-of-the-Art Methods
We conduct a comprehensive evaluation of our proposed MobileGeo framework against a wide array of recent state-of-the-art methods. As presented in Table I, the comparison focuses on both retrieval accuracy (Recall@1, AP) and model efficiency (Parameters, FLOPs, FPS).
Superior Efficiency. Our core model, MobileGeo, exhibits high computational efficiency. With only 28.57M parameters and an exceptionally low 4.45G FLOPs, it is by far the most lightweight and computationally inexpensive model among all compared methods. To put this in perspective, compared to the recent efficient model MEAN [4], MobileGeo reduces FLOPs by a remarkable factor of 5.8 and parameters by 21.2%. This optimization directly translates to a massive 3.3 increase in FPS, reaching 1022 images/second, a critical capability for real-time deployment on resource-constrained platforms.
State-of-the-Art Accuracy. Building upon this highly efficient foundation, our full model, MobileGeo†, incorporates the MSRM as a post-processing step achieves an impressive 97.15% Recall@1 and 97.50% AP in the primary DroneSatellite retrieval task. This represents a substantial absolute improvement of 3.30% in R@1 over our efficient baseline and significantly surpasses the previous best-performing method, DAC [10], all while operating with over 20 fewer FLOPs (4.45G vs. 90.24G).
@@ IV-C Unsupervised Domain Adaptation Results
As demonstrated in Table II, MobileGeo exhibits exceptional generalization in the DroneSatellite task. While MEAN [4] shows strong performance at 150m, our model consistently outperforms all state-of-the-art methods at 200m, 250m, and 300m. Notably, at the 300m range, MobileGeo achieves a top-ranking R@1 score of 97.87, surpassing the next-best competitor by a significant 3.24 percentage points. It is crucial to highlight that MobileGeo utilizes only 4.45G FLOPs, compared to the 26.18G FLOPs required by MEAN.
The superior generalization of MobileGeo is more pronounced in the SatelliteDrone task, as shown in Table III. In this scenario, our model unequivocally achieves the best performance across all evaluation altitudes. For instance, it surpasses the strong baseline DAC [10] by 3.75 percentage points in R@1 at the 250m altitude.
IV-D Multi-weather drone imagery degradation Results
As shown in Figure 6, to evaluate the model’s robustness in practical scenarios, we conducted extensive experiments on the University-1652 dataset under various environmental degradations, including fog, rain, darkness, and wind. As presented in Table IV, these tests assess the model’s ability to maintain high accuracy even when the visual quality of drone imagery is severely compromised.
In the DroneSatellite retrieval task, our proposed MobileGeo demonstrates remarkable resilience and establishes a new state-of-the-art across all tested conditions. From normal weather to the most severe degradations like darkness and combined fog with rain, MobileGeo consistently achieves the highest Recall@1 and AP scores. For instance, under dark conditions, it outperforms the next-best method, MEAN [4], by a substantial margin of 5.37 percentage points in R@1 (93.27% vs. 87.90%).
In the more challenging SatelliteDrone task, MobileGeo continues to show highly competitive performance. In the most complex scenarios such as ”Fog+Rain” and ”Wind”, MobileGeo takes the lead in both metrics, achieving the best R@1 and AP scores. This proves that even when satellite images must be matched against heavily distorted drone views, our model’s feature representation remains robust and effective.
@@ IV-D Multi-weather drone imagery degradation Results
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 70.05 | 74.93 | 80.68 | 83.90 | 87.35 | 89.72 | 90.03 | 91.91 |
| DAC[10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 76.65 | 80.56 | 86.45 | 89.00 | 92.95 | 94.18 | 94.53 | 95.45 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 81.73 | 87.72 | 89.00 | 91.00 | 92.13 | 93.60 | 94.63 | 95.76 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 79.85 | 83.32 | 91.10 | 92.79 | 95.57 | 96.45 | 97.87 | 98.32 |
@@
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 83.75 | 73.83 | 91.25 | 83.42 | 93.75 | 89.07 | 93.75 | 90.66 |
| DAC[10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 87.50 | 79.87 | 96.25 | 88.98 | 95.00 | 92.81 | 96.25 | 94.00 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 91.25 | 81.50 | 96.25 | 89.55 | 95.00 | 92.36 | 96.25 | 94.32 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 92.50 | 83.81 | 97.50 | 91.75 | 98.75 | 94.59 | 97.50 | 96.04 |
@@
| MobileGeo (Ours) | 4.45 G | 93.27 | 94.34 | 95.44 | 89.95 |
| (g) Wind | |||||
| LPN [2] | 65.39 G | 66.46 | 70.35 | 84.14 | 67.35 |
| MuSeNet [53] | 42.37 G | 69.45 | 73.22 | 86.31 | 70.03 |
| Sample4Geo[14] | 90.24 G | 83.39 | 89.51 | 95.29 | 87.06 |
| MEAN[4] | 26.18 G | 89.27 | 91.01 | 95.44 | 86.05 |
| MobileGeo (Ours) | 4.45 G | 93.45 | 94.48 | 95.58 | 91.61 |
@@ IV-E Anti-offset Generalization Results
In real-world drone-view geo-localization, the captured query image is often not perfectly centered or aligned with its corresponding satellite-view gallery image. This spatial misalignment can be caused by variations in camera angle or differences in viewpoint. A robust model must be able to generalize well despite such spatial offsets.
@@
The comprehensive results are presented in Table V. The findings demonstrate the superior robustness of our proposed MobileGeo model. For example, under the most severe shift of pixels, where other models experience a significant performance collapse, MobileGeo maintains an exceptional Rank-1 accuracy of 93.86%. This represents a massive improvement of +16.92% over the second-best method, SDPL.
It is crucial to note that while the performance of all models degrades as the offset magnitude increases, MobileGeo’s degradation is significantly more graceful.
| Padding Pixel | FSRA (98.05 G) | LPN (65.39 G) | SDPL (69.71 G) | MobileGeo (4.45 G) | ||||
@@
| HD-CVGL | MSRM | DroneSatellite | SatelliteDrone | ||||
@@ IV-F Ablation Studies
Our MobileGeo framework is architected training-time optimization and lightweight inference-time post-process. To validate the specific contributions, we conducted extensive ablation studies, with the results summarized in Table VI.
@@
Effectiveness of the Training-Phase Framework. The core objective of our training strategy is to distill rich, view-invariant knowledge into a single, efficient network, avoiding the need for multi-branch architectures during inference. Our analysis starts with a baseline model (row 1) achieving 86.44% Recall@1. By introducing Self-Distillation (SD), performance improves significantly to 91.46%, which ensures that the model can fully leverage its own multi-level feature hierarchy for more robust representation learning. Following this, integrating the UAFA module further advances the Recall@1 to 91.93%. Finally, Cross-Distillation (CD) elevates performance to 93.87%, completing the HD-CVGL framework. This step-wise improvement validates our training-phase design, which produces a powerful feature extractor ready for efficient inference. As visualized in Figure 8 and Figure 9, our MobileGeo model learns a more discriminative feature distribution compared to the baseline. The features produced by MobileGeo exhibit smaller intra-class variance and larger inter-class distance, indicating a well-structured feature space.
@@
Effectiveness of the Inference-Phase Module. After establishing a strong student model through our training strategy, we evaluate the contribution of the MSRM. The final row of Table VI shows the impact of applying MSRM to the descriptors generated by our fully trained model. The result is a remarkable jump in performance to 97.15% Recall@1. This +3.28% gain over the already powerful base model demonstrates that by refining the feature set at inference time, MSRM significantly enhances localization precision, fulfilling the ”Precision-Focused” aspect of our MobileGeo paradigm without altering the underlying network architecture. Figure 10 illustrates the performance of different selected views. The model achieves its peak performance with 40 views. Even at its lowest setting (20 views), our approach still significantly outperforms the baseline without the MSRM module. As illustrated by the top-5 retrieval results on the University-1652 and SUES-200 datasets (Figure 11 and Figure 12), MobileGeo demonstrates strong generalization and robustness in cross-view image matching capabilities, enabling accurate localization even in GPS-denied outdoor environments.
@@
Edge Device Deployment. To demonstrate the practical applicability of our approach, as shown in Figure 14, we evaluate MobileGeo on two representative edge platforms: NVIDIA Jetson TX2 and AGX Orin. As shown in Figure 13, our method achieves exceptional efficiency with only 4.45 GFLOPs, representing a 95.1% reduction compared to DAC. This dramatic reduction translates directly to superior real-time performance: MobileGeo achieves 27.1 FPS on the resource-constrained TX2 (7.7 faster than DAC’s 3.5 FPS) and an impressive 251.5 FPS on AGX Orin (6.1 faster than DAC). Furthermore, even when compared to the efficient MEAN method, our approach exhibits substantial speedups. On the TX2, MobileGeo is 2.9 times faster, and on the AGX Orin, it achieves a 2.8 speedup. These results underscore the exceptional performance of MobileGeo in resource-constrained environments, highlighting its potential for real-world applications on mobile and edge devices.
As mentioned above, the efficiency gains observed are not only a result of reduced computational overhead but also reflect the method’s ability to optimize resource usage effectively, a critical factor in edge computing.
@@ IV-G Limitations and Future Works
Despite its promising performance, MobileGeo has several limitations that guide our future work. The current model, while excelling at DroneSatellite task exhibits an asymmetric performance due to its focus on the multi-view refinement of drone features. Future efforts will explore the intra-relationships within the satellite data to build a more generalized cross-view feature space. Additionally, MobileGeo’s lightweight design faces challenges with extreme low-altitude distortions, which could be mitigated by developing adaptive, scale-aware mechanisms or pre-transformation techniques. To bridge the gap between efficient image-level retrieval and the need for precise localization, our ultimate goal is to evolve beyond static images. We plan to leverage multiview sequence data and develop an end-to-end framework that integrates coarse scene identification with fine-grained pose estimation. This would enable robust and accurate geo-localization within large-scale satellite imagery database.
To further enhance the real-time capabilities of MobileGeo, we aim to incorporate dynamic updates directly on mobile devices. By leveraging the capabilities of 5G networks and cloud-edge collaboration, we plan to optimize the model through distributed training, enabling seamless real-time geo-localization even in challenging unseen environments. This strategy will ensure that MobileGeo can not only handle large-scale datasets but also adapt quickly to changing geographical conditions, providing both efficiency and precision in mobile and drone-based applications.
@@ V Conclusion
In this paper, we introduced MobileGeo, a mobile-friendly framework that follows a precision-focused efficient design paradigm. Our core strategy shifts computational complexity to the training stage, enabling superior performance on resource-constrained devices. We achieve this through two innovations: 1) A comprehensive Hierarchical Distillation (HD-CVGL) strategy during training, which incorporates our Uncertainty-Aware Prediction Alignment (UAPA) to robustly handle data imbalance and domain discrepancies, producing a highly discriminative yet compact student network without any inference overhead. 2) A lightweight Multi-view Selection Refinement Module (MSRM) at inference, which uses mutual information to select and fuse the most informative views, boosting accuracy while minimizing feature matching cost. Extensive experiments demonstrate that MobileGeo outperforms previous state-of-the-art methods, e.g., achieving improvements of 4.19% in AP on University-1652 while being over 5 efficient in FLOPs and 3 faster during inference. Crucially, its 251.5 FPS on AGX Orin validates its practicality for outdoor drone geo-localization in GNSS-denied environments on resource-constrained mobile and edge devices.